OpenStudy (pphalke):

$8000 is invested in a simple annual interest rate of 6% find the balance of the investment after 5 years i got $10,400 and it was a wrong answer can i get help

OpenStudy (campbell_st):

well the simple interest formula is \[I = P \times \frac{r}{100} \times t\] ypu have P = 8000 r = 6 and t = 5 so calculate the interest the the Balance or future value is Future Value = Principal + Interest

OpenStudy (mathstudent55):

I don't understand why your answer is wrong.

OpenStudy (pphalke):

i got 24

OpenStudy (pphalke):

I=8000*0.06/100*5

OpenStudy (pphalke):

would it be 24,000??

OpenStudy (pphalke):

sorry its 7976 right @campbell_st

OpenStudy (mathstudent55):

What is 6% of $8000?

OpenStudy (campbell_st):

as I said in the posting r = 6 so you need \[I = 8000 \times \frac{6}{100} \times 5\]

OpenStudy (mathstudent55):

$7976 is less than $8000. You start with $8000, earn interest for 5 years, and end up with less than $8000? How can that be?

OpenStudy (campbell_st):

when you get the value of the interest add it to the principal.

OpenStudy (pphalke):

its 5600

OpenStudy (pphalke):

right @campbell_st

OpenStudy (mathstudent55):

Are you asking if 5600 is the interest earned in 5 years?

OpenStudy (campbell_st):

is that the interest calculation..?

OpenStudy (pphalke):

yes itd what i got

OpenStudy (mathstudent55):

You wrote above in your post that you got $10,400 for the answer. How did you get that? Did you calculate it? How did you calculate $10,400?

OpenStudy (pphalke):

i dont know but i know it was wrong

OpenStudy (mathstudent55):

Where did you get the idea it's wrong?

OpenStudy (pphalke):

when i submitted the answers it came wrong, the professor marked it wrong

OpenStudy (mathstudent55):

Also, if you are finding an interest rate, then can you tell me what 8% simple interest on $8000 for 1 year is? Simply put, what is 6% of $8000?

OpenStudy (mathstudent55):

To find a percent of a number, multiply the percent by the number. You change the percent to a decimal by moving the decimal point two places to the left. Then you multiply by the number.

OpenStudy (mathstudent55):